in a chemistry class, 7 litters of a 4% silver iodide solution must be mixed with a 10% solution to get a 6% solution. how many liters of the 10% solution are needed?
Answer by stanbon(75887)

in a chemistry class, 7 liters of a 4% silver iodide solution must be mixed with a 10% solution to get a 6% solution. how many liters of the 10% solution are needed?
---------------------------------
Equation:
silver + silver = silver
0.04*7 + 0.10x = 0.06(7+x)
------------------------------------
Multiply thru by 100 to get:
4*7 + 10x = 6*7 + 6x
4x = 2*7
x = 3.5 liters (amt. of 10% solution needed)
=================================================
Cheers,
Stan H.